# LLM-Human Pipeline for Cultural Context Grounding of Conversations

### v1

@@ Abstract
Conversations often adhere to well-understood social norms that vary across cultures. For example, while addressing parents by name is commonplace in the West, it is rare in most Asian cultures. Adherence or violation of such norms often dictates the tenor of conversations. Humans are able to navigate social situations requiring cultural awareness quite adeptly. However, it is a hard task for NLP models.
In this paper, we tackle this problem by introducing a Cultural Context Schema for conversations. It comprises (1) conversational information such as emotions, dialogue acts, etc., and (2) cultural information such as social norms, violations, etc. We generate $\sim$110k social norm and violation descriptions for $\sim$23k conversations from Chinese culture using LLMs. We refine them using automated verification strategies which are evaluated against culturally aware human judgements. We organize these descriptions into meaningful structures we call Norm Concepts, using an interactive human-in-loop framework. We ground the norm concepts and the descriptions in conversations using symbolic annotation. Finally, we use the obtained dataset for downstream tasks such as emotion, sentiment, and dialogue act detection. We show that it significantly improves the empirical performance.
LLM-Human Pipeline for Cultural Context Grounding of Conversations
@@ 1 Introduction
We leverage LLMs such as GPT-$3.5$ to generate a large but potentially noisy corpus of conversation-specific cultural information. Then, we leverage an interactive human-in-loop process (Alg. 1) that efficiently utilizes culturally proficient human annotation to organize this information into meaningful concepts. We further ground the generated descriptions in the conversation details, such as participants, their relationships, and their behavior using symbolic annotation. Following this, we experiment with automated verification strategies to filter the obtained cultural information at scale. Then, we organize the conversations and the obtained cultural information into a meaningful schema structure. Finally, we propose a neural graph schema model that leverages obtained data to improve the empirical performance on conversation understanding tasks such as emotion detection, sentiment detection, and dialogue act detection across multiple datasets. Overall, our contributions can be summarized as follows:
@@
    We introduce Norm Concepts and employ a human-in-loop (HiL) framework to create human-validated concepts supported by data.
@@
    We leverage the cultural schema information to improve downstream conversational task performance. We present meaningful visualizations and human evaluation experiments to showcase the quality. [^1] We will release our data and code under MIT license
@@ 2 Related Work
Social Norm Datasets: Few recent works have leveraged LLMs such as GPT-3 to create useful cultural norm datasets using structure prompting approaches (Fung et al., 2023; Li et al., 2023; Ziems et al., 2023). While these works either focus on a small set or synthetically generated conversations, we generate $63k$ norm descriptions for $23.5k$ real conversations and ground them using a principled cultural grounding pipeline.
@@ 3 Data
(b) Example norm concepts and symbolic explanations
Figure 1: Proposed Cultural Context Schema Structure for Conversations
@@
LDC CCU ZH Text Dataset: This dataset is a conversational understanding dataset from the DARPA CCU program. This dataset consists of text, audio, and video data in several languages. We use the text portion of the Chinese dataset for our experiments. It contains annotations for emotion, dialogue act, social norm violation status, and conversation change points. We focus on emotion and dialogue act identification tasks. This dataset also has metadata about settings, age of participants, familiarity, etc, annotated.
@@ 4 Cultural Context Grounding
    Formalize relevant cultural context for better conversational understanding
@@
    Evaluate the obtained cultural context dataset for correctness and usefulness
@@ 4.2.4 Automated Verification
Figure 3: Automated Verification Flowcharts. S* maps to pipeline stages in Fig. 2

### v2

# LLM-Human Pipeline for Cultural Grounding of Conversations

@@ Abstract
Conversations often adhere to well-understood social norms that vary across cultures. For example, while addressing work superiors by their first name is commonplace in the Western culture, it is rare in Asian cultures. Adherence or violation of such norms often dictates the tenor of conversations. Humans are able to navigate social situations requiring cultural awareness quite adeptly. However, it is a hard task for NLP models.
In this paper, we tackle this problem by introducing a Cultural Context Schema for conversations. It comprises (1) conversational information such as emotions, dialogue acts, etc., and (2) cultural information such as social norms, violations, etc. We generate $\sim$110k social norm and violation descriptions for $\sim$23k conversations from Chinese culture using LLMs. We refine them using automated verification strategies which are evaluated against culturally aware human judgements. We organize these descriptions into meaningful structures we call Norm Concepts, using an interactive human-in-the-loop framework. We ground the norm concepts and the descriptions in conversations using symbolic annotation. Finally, we use the obtained dataset for downstream tasks such as emotion, sentiment, and dialogue act detection. We show that it significantly improves the empirical performance.
LLM-Human Pipeline for Cultural Grounding of Conversations
@@ 1 Introduction
We leverage LLMs such as GPT-$3.5$ to generate a large but potentially noisy corpus of conversation-specific cultural information. Then, we leverage an interactive human-in-the-loop process (Alg. 1) that efficiently utilizes culturally proficient human annotation to organize this information into meaningful concepts. We further ground the generated descriptions in the conversation details, such as participants, their relationships, and their behavior using symbolic annotation. Following this, we experiment with automated verification strategies to filter the obtained cultural information at scale. Then, we organize the conversations and the obtained cultural information into a meaningful schema structure. Finally, we propose a neural graph schema model that leverages obtained data to improve the empirical performance on conversation understanding tasks such as emotion detection, sentiment detection, and dialogue act detection across multiple datasets. Overall, our contributions can be summarized as follows:
@@
    We introduce Norm Concepts and employ a human-in-the-loop (HiL) framework to create human-validated concepts supported by data.
@@
    We leverage the cultural schema information to improve downstream conversational task performance. We present meaningful visualizations and human evaluation experiments to showcase the quality. [^1] Code and Data: https://github.com/pujari-rajkumar/cultural-schema-naacl2025
@@ 2 Related Work
Social Norm Datasets: A few recent works have leveraged LLMs such as GPT-3 to create useful cultural norm datasets using structure prompting approaches (Fung et al., 2023; Li et al., 2023; Ziems et al., 2023). While these works either focus on a small set or synthetically generated conversations, we generate $63k$ norm descriptions for $23.5k$ real conversations and ground them using a principled cultural grounding pipeline.
@@ 3 Data
(b) Norm Concepts & Symbolic Explanations example. $S_{*}$ denotes relevant pipeline stages from Fig. 2.
Figure 1: Proposed Cultural Context Schema Structure for Conversations with an example instance
@@
LDC CCU ZH Text Dataset: This dataset is a conversational understanding dataset from the DARPA CCU program. This dataset consists of text, audio, and video data in several languages. We use the text portion of the Chinese dataset for our experiments. It contains annotations for emotion, dialogue act, social norm violation status, and conversation change points. We focus on emotion and dialogue act identification tasks. This dataset also has metadata about settings, age of participants, familiarity, etc, annotated. The dialogue act annotation for this dataset is not complete. The annotation is focused only on a pre-determined set of dialogue acts. We deal with this by using other class for unlabeled data.
@@ 4 Cultural Context Grounding
    Formalize relevant cultural context for better conversational understanding.
@@
    Evaluate the obtained cultural context dataset for correctness and usefulness.
@@ 4.2.4 Automated Verification
Figure 3: Automated Verification Flowcharts. $S_{*}$ denotes pipeline stages from Fig. 2
@@
4.3 Comparison with Existing Norm Datasets
Three of the prominent existing works that address social norms in conversation are NORMSAGE (Fung et al., 2023), NORMDIAL (Li et al., 2023) and NORMBANK (Ziems et al., 2023).
NORMSAGE presents a framework to obtain norms using LLMs and then perform self-verification. Their main contribution is the pipeline rather than a dataset. In contrast, we gather norm descriptions, operationalize a structured organization and filtering pipeline that efficiently uses and amplifies culturally proficient human annotation.
NORMDIAL is closest to our work as they propose a human-in-the-loop approach with LLM to obtain situation-specific norms. However, we introduce norm concepts that aggregate situation-specific descriptions into more general behavioral concepts. We also present robust automated verification, which is evaluated using multi-point human evaluation. The scale of our dataset is larger.
NORMBANK uses situational attributes that align with the factual segment of our schema. However, unlike our work, these norms are not derived from real conversations but are instead defined by a list of high-level scenarios. Our dataset focuses on capturing nuanced human behavior within a specific culture, whereas NORMBANK aims for broader, more general cultural coverage.
@@ 8 Conclusion
Acknowledgments
We thank the reviewers for their insightful comments that helped improve the paper. This work was supported by NSF CAREER award IIS-2048001 and the DARPA CCU program. The contents are those of the author(s) and do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
@@ Acknowledgments
LLM-generated cultural norms could potentially perpetuate stereotypes, contain hallucinations, bias, or other harmful distributional characteristics that could be challenging to detect and filter. In this work, we address these issues such as hallucinations, and stereotypes by (1) involving culturally proficient humans in the norm concept creation stage and (2) also performing principled automated filtering which is again evaluated against human annotation. In our pipeline, humans are not just aggregating the descriptions that they see, but rather they are using their cultural expertise to identify norm concepts that are prominent in their culture and are also supported by data. Despite these steps, the output is not necessarily completely absolved of these issues. On the other hand, human-generated datasets have also been shown to contain harmful biases (Hovy and Prabhumoye, 2021; Geva et al., 2019; Gautam and Srinath, 2024; Davidson et al., 2019; Das et al., 2024; Doughman and Khreich, 2022). LLMs are powerful NLP tools that allow us to explore tasks that were not practical before their advent. A case in point is our paper. It would have been highly impractical or hugely expensive to collect culturally proficient human annotations for cultural norms at this scale without LLMs. We believe that a pragmatic approach to this issue would be to carefully document and evaluate biases at each stage of the pipeline and invest community effort in building de-biasing techniques, bias-free training approaches, and so on. This could potentially lead us to creating fair(er) systems. We believe that a pragmatic approach of learning to build better guard rails, mitigating the harmful effects, and being aware of these biases could be a more fruitful path forward.